A Complete Bibliography of Publications in Synthese, 1930–1939

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

-Theorie [219, 75]. -Theory [219].

14th [253]. 1938 [253]. 1939 [339, 342]. 19th [253].

28 [342].

A. [246, 301, 300]. Aan [74]. Abroad [18, 22, 34, 41, 48, 54, 61, 68, 88, 105, 170]. Absolute [235, 235]. against [124, 80]. Age [268]. Alfred [300]. Algemeen [288]. algemeene [59]. Altruism [348]. Altruisme [348]. Amsterdam [313]. Analizo [145]. Analyse [278, 91]. Analysis [29, 145, 91, 278, 290]. Angst [111]. animal [24, 39, 53]. Antithesen [135]. Antithesis [135]. Anxiety [111]. Any [228]. Apologetica [182]. Apologetics [182]. Apologie [16]. Apology [16]. appeared [342]. April [362]. Apriori [115, 122, 128]. argeloosheid [80]. Aristoleles [11, 15]. Aristoteles [3]. Aristotle [3, 11, 15]. Arrived [179].

Art [162, 59]. articles [356]. Astapowo [85]. Astonishing [111]. Athene [195]. Athens [195]. attaining [296]. August [256]. Autonomie [319, 309]. Autonomy [319, 309].

Background [188, 181]. Be [129]. Beauty [116, 30]. beeld [36]. beeldspraak [2, 9]. Beginselen [230]. Begriffe [191]. Begriffsbestimmung [71]. Begrip [91, 66]. Begrippen [276]. Begripsontleding [29]. Begripsvorming [294]. Behoefte [139]. Being [338, 137]. Belang [332]. Beleefd [129]. Beleving [23, 103]. Belevingswaarheid [82]. Believe [204]. Bemerkungen [191]. Beoordeeling [282]. Berdjajew [151]. Bergson [268, 181]. Bergsonian [188]. Bergsonienne [188]. Beschouwingen [162, 132, 96, 104, 110, 133, 140]. Best [297]. Beteekenis [276, 81, 77]. Beteekent [204]. betekenis [29]. Betrekkingen [285]. Between [130, 97]. Bewustzijn [146]. Bij [311, 297, 131, 24, 39, 53]. Bijdrage [233, 91]. Biographical [298]. Biologie [341, 216, 340]. Biologist [114]. Biology [341, 340, 216]. Bioloog [114]. Board [313]. Boekbespreking [26, 43, 89, 193, 201, 244, 251, 279, 272, 171, 210]. Boeken [179, 236, 273, 327]. Book [26, 43, 89, 193, 201, 244, 251, 279, 272, 171, 210]. Books [356, 179, 236, 273, 327]. Bouw [38, 47, 31]. Bouwen [97]. Bovenpersoonlijke [152, 163]. Bridge [130]. Brief [64]. Brug [130]. Bruhl [281]. Buber [63]. Buddhism [349]. Buddhisme [349]. Buitenland [18, 22, 34, 41, 48, 54, 61, 68, 88, 105, 170]. Buitenlandsche [243, 252, 274, 280, 291, 299, 306, 312, 318, 326, 336, 345]. Bussy [347].

call [342]. Cambridge [253]. Captures [3]. Case [136, 136]. Category [3, 11, 15]. Catholic [292, 81, 182]. Catholicism [118]. Causal [276]. Causale [276]. Causaliteit [276]. Causaliteitsbeginsel [4]. Causality [131, 4, 276, 46]. Causerie [339]. Chapters [184, 206, 234, 317]. Chestov [84, 27, 245]. Chineezen [224]. Chinese [224]. Christendom [230]. Christianisme [231]. Christianity [230, 231]. Cities [158]. clearing [290]. Cognitive [220]. collective [353]. Collectivism [352]. Collectivisme [352]. College [253]. Comments [154, 228]. Compared [3]. Complaint [160]. Concept [91, 191]. Conceptie [197]. Conception [197]. concepts [276]. Conceptualization [294]. Conference [339, 328]. Confidences [83]. Congress [253, 361]. Consciousness [146]. Considerations [132, 96, 104, 110, 133, 140]. Constant [82]. Constateerwaarheid [82]. Construction [97, 38, 47, 31]. Contemporary [225, 32, 8, 17, 81, 57, 265]. Contra [111]. Contrast [111]. Contribution [233, 91]. Correspondence [12, 19, 33, 42, 49, 55, 62, 69, 79, 202]. Correspondentie [12, 19, 33, 42, 49, 55, 62, 69, 79, 202]. Cosmogonie [96, 104, 110, 200, 200]. Cosmogony [96, 104, 110]. Cosmos [311]. Creator [155, 156]. Criminologie [184, 206, 234, 317]. Criminology [184, 206, 234, 317]. Crisis [320, 83, 221, 322, 321, 222, 4]. Criticism [315]. Critiek [315]. Cult [198]. Culte [198]. Culture [265, 91, 20, 58]. Culturiseeren [20]. Cultuur [91, 58]. Cultuurphilosophie [265]. Current [238].

December [261]. Declaration [168]. definition [71]. Degradation [228]. Demand [226]. Demanding [21]. Demoeizame [21]. Denken [263, 176, 315, 213, 16, 267, 147, 74, 21, 138]. denkleven [292, 81]. Deutung [214]. Development [180, 287]. Deze [276]. Dezen [139]. Dialogen [109, 125, 134]. Dialogue [109, 125, 134]. d’Idées [295]. Dier [24, 39, 53]. difference [97]. dimensie [141]. Dimension [141]. Ding [189, 196, 215]. Diseases [168]. Divine [186]. Doctrine [224]. Doeltreffende [270]. does [204]. door [339]. Doorgangsstadia [158]. Dostoevsky [344, 343]. Dostojewsky [344, 343]. Dr [249]. Draw [136]. Dream [137]. Drie [169]. Drieëenheid [102]. Driesch [337, 114]. Droom [137]. Dust [205]. Dutch [26, 43, 12, 19, 33, 42, 49, 55, 62, 69, 79, 23, 57, 1, 64, 50, 18, 22, 34, 41, 48, 54, 61, 68, 89, 162, 132, 154, 141, 142, 144, 88, 105, 170, 160, 116, 129, 233, 243, 252, 193, 201, 244, 251, 202, 211, 179, 212, 177, 185, 207, 247, 187, 248, 194, 203, 236, 238, 218, 178, 239, 279, 274, 280, 294, 291, 299, 306, 312, 318, 326, 336, 345, 339, 320, 277, 307, 346, 265, 337, 313, 269, 285, 329, 264, 273, 327, 262]. Dutch [302, 284, 292, 331, 93, 224, 152, 163, 343, 319, 332, 347, 314, 263, 275, 282, 109, 125, 134, 268, 197, 176, 235, 311, 199, 28, 85, 195, 180, 139, 250, 186, 289, 288, 169, 315, 213, 59, 149, 97, 124, 270, 225, 349, 32, 8, 17, 328, 272, 16, 31, 38, 47, 4, 83, 21, 72, 52, 96, 104, 110, 158, 130, 200, 267, 230, 67, 119, 147, 241, 286, 155, 135, 184, 206, 234, 317, 217, 117, 36, 133, 140, 13, 114, 94, 99, 106, 136, 276, 341, 190, 25, 5, 159, 98, 75, 24, 39, 53]. Dutch [6, 87, 112, 148, 216, 240, 303, 3, 11, 15, 29, 7, 171, 210, 228, 227, 324, 350, 102, 219, 174, 258, 80, 181, 352, 334, 333, 221, 300, 297, 229, 157, 81, 322, 91, 108, 143, 209, 74, 82, 60, 14, 46, 20, 30, 51, 37, 66, 2, 9, 146, 103, 131, 165, 95, 111, 123, 90, 118, 86, 182, 204, 242, 342, 138, 168, 323, 335, 309, 77, 58, 137, 348, 73, 283, 205, 342]. Dynamics [186]. Dynamiek [186].

Economic [284, 286, 287]. Economie [25, 5, 159, 98]. Economische [284, 286]. Economy [25, 5, 159, 98]. Eduard [45, 238, 239]. Education [178]. eener [59]. Eenheidswetenschap [174]. Eenige [154]. Eenzame [165]. Eerste [339, 328]. Effective [270]. effects [287]. Effusione [127, 121, 126]. Egoism [348]. Egoisme [348]. Einheitswissenschaft [175, 359]. Einige [191]. Einzelne [65]. Einzige [65]. Emotie [131]. Emotief [267]. Emotion [131]. Emotional [267]. Empfindungen [278]. Empire [73]. Empirical [169, 174, 175, 107]. Empiricism [360]. Empirie [73]. Empirische [169]. Empirischen [107]. Empirismus [360]. Empiristische [174, 175]. Encyclopedia [359]. England [253]. Enige [228]. enkele [32, 8, 17]. Entelechie [307, 307, 308]. Entelechy [308]. Enzyklopaedie [359]. Erfolg [156]. Erkenntnistheorie [220]. Erlebnis [122, 115, 128]. Ernst [278, 277, 278]. Errata [56, 78, 113, 161, 223, 237]. Erratum [40, 100, 120, 150]. Esperanto [145]. Essence [52, 118, 143]. Ethical [347]. Ethische [347]. Ethnologie [225]. Ethnology [225]. Étude [271]. Evolutie [225]. Evolutie-Gedachte [225]. Evolution [225]. Evolution-Thought [225]. Existence [117]. Existentie [117]. Existentie-Philosophie [117]. Experience [23, 122, 103, 115, 128]. Extraordinary [152, 153].

Failure [156]. Faith [190, 131]. February [357]. Fifth [361]. First [339, 328]. Fledderus [304]. folk [93]. following [313, 342]. follows [247]. Fond [188]. Forderung [226]. Foreign [243, 252, 274, 280, 291, 299, 306, 312, 318, 326, 336, 345]. Formation [295, 295]. Forum [254, 256, 257, 259, 260, 261, 354, 357, 358, 362, 255]. Fourth [141, 253]. Freedom [343, 94, 99, 106, 344, 107, 275]. Freienfels [101]. Freiheit [107, 344]. French [295, 271, 198, 293, 188, 231]. Friedrich [211, 157]. Functie [297]. Function [297].

G [248, 249, 339, 264]. G. [248, 264]. Game [199]. Gang [346, 21]. Gebaar [60]. Gedachte [18, 22, 34, 41, 48, 54, 61, 68, 88, 105, 225]. Gedachten [170, 165]. Gedeelte [288]. Geest [205]. Geesteswetenschap [59]. Gehouden [339]. Geistige [226]. Geloof [190]. Gelooven [204]. gene [74]. Geneeskunde [250, 289]. General [288, 59]. Genie [154]. Genius [154]. German [278, 330, 153, 164, 344, 266, 65, 214, 115, 122, 128, 189, 196, 215, 308, 338, 355, 10, 156, 107, 340, 360, 71, 226, 351, 220, 175, 258, 359, 353, 321, 208, 232, 191, 316, 325]. Gesamtgeschichte [266]. Geschichtliche [226]. Geschichtswissenschaft [266]. Geschriften [179, 236, 273, 327]. Gestalt [75, 75]. gesture [60]. Getallen [95]. Getting [297]. Gewaarwording [23]. Geweldmiddel [149]. Girton [253]. Gisteren [7]. Golden [268]. Gouden [268]. Groep [242]. grond [72]. ground [72]. Group [185, 271, 269, 227, 212, 177, 242]. Groupe [271]. Gunstigste [297].

H [35, 246]. haar [93]. Hans [337, 114]. Hartmann [45]. Hartog [246]. Hedendaagsche [57, 265, 32, 8, 17, 81]. Heilige [95]. Held [339]. Heteronomie [319]. Heteronomy [319]. Heymans [248, 249, 264]. Hildebrandt [214, 213]. Hildebrandts [214]. Hindhu [3, 11, 15]. Hindhu-wijsbegeerte [3, 11, 15]. Hindu [3, 11, 15]. Historical [226]. Historie [87, 112, 148]. History [162, 87, 112, 148, 266, 59]. Hollaendischer [258]. Holy [95]. Hoofddirectie [313]. Hoofdstukken [184, 206, 234, 317]. Huidige [238]. Human [163, 164, 152, 153]. humanities [59].

I. [347, 189]. I.S.S [269, 227]. I.S.S.G [269]. Ideal [267]. Ideas [295]. Ideatief [267]. Iets [224]. II [338]. II. [338]. III [141]. image [36, 2, 9]. Importance [332]. impotence [80]. Indifference [207]. individual [65]. Individualism [352]. Individualisme [352]. Industrial [285]. Industrieele [285]. Influences [286]. Information [123, 123]. Ingekomen [179]. Inleiding [1, 142, 144, 199]. Inlichting [123]. Institute [173, 183, 285]. Instituut [285]. International [253, 361, 271, 285, 359, 173, 183, 212, 177, 185, 271, 269, 227]. Internationale [185, 285, 227, 359, 212, 177, 269]. Interpretation [214]. Introduction [1, 142, 144, 199]. Intuïtieve [67]. Intuitive [67]. Invloeden [286].

J [347]. Jaar [268]. James [35]. January [354]. Jasnaja [85]. Jasnaja-Poljana [85]. Jeans [35]. Jerusalem [195]. Jeruzalem [195]. Journal [342]. Judgment [282]. Juli [339]. July [253, 339, 254]. Juridical [342]. Juridische [241]. Juristenblad [342].

Kali [198, 197]. Kant [3, 11, 15, 166, 167]. Kategorieënleer [3, 11, 15]. Katholicisme [118]. Katholieke [292, 81, 182]. Kenbaarheid [207]. Kennis [219]. Kennis-Theorie [219]. Klacht [160]. Knowability [207]. knowledge [67, 219]. Köhler [75]. Kollektivlebens [353]. Korte [64]. Kosmos [311]. Krachtproductie [286]. Krise [321]. Kunstgeschiedenis [162, 59]. Kurt [214, 213].

L [304]. Langage [295]. Language [294, 295, 228, 270]. Learning [3, 11, 15]. Leer [224]. Legal [241]. Leo [84, 27, 245]. Leuzen [149]. Leven [13, 276, 303]. Levens [333]. Levens- [333]. Levenspad [349]. Levenspsychologie [218]. Levensverschijnselen [13]. Levy [281]. Levy-Bruhl [281]. Liberal [360]. Liberaler [360]. licht [59]. Life [218, 13, 303, 333, 292, 349, 276, 353, 81, 13, 303]. light [59]. Lighting [123]. Lijn [247]. Line [247]. Linguistic [271, 332]. Linguistique [271]. logic [222, 355, 351]. Logica [221]. Logik [355, 351]. Logistic [321]. Logistics [158]. Logistiek [158]. Logistischen [321]. Lonely [165]. Lucien [281].

Maatschappij [242]. Mach [278, 277, 278]. Magazines [243, 252, 274, 280, 291, 299, 306, 312, 318, 326, 336, 345]. Main [313]. Man [331, 130, 155, 24, 39, 53, 330, 334]. Manifestatie [235]. Manifestation [235]. mankind [266]. Mannoury [339]. March [358]. Martin [63]. Mary [304]. Mass [331, 329, 330]. Massa [331]. Massapsychologie [329]. Masse [330]. Matematika [145]. Materie [38, 47, 52, 31]. Mathematical [145]. Mathematics [116, 72]. Matter [38, 47, 52, 31]. mean [204]. Meaning [208, 276, 29, 81]. means [149]. measure [141]. Mechanism [6]. Mechanisme [6]. Mededeelingen [194, 203]. Mediations [194, 203]. Medicine [250, 289, 290]. Meetings [190]. Men [311]. Mensch [331, 330, 311, 130, 155, 24, 39, 53, 334]. Menschelijke [152, 163]. Menschenkennis [67]. Menschheit [266]. Menschliche [153, 164]. Mère [198]. Metafysika [158]. Metaphysica [187, 248, 264, 73]. metaphysical [72, 10]. Metaphysics [187, 248, 264, 73, 158]. metaphysische [72, 10]. meten [141]. Method [270]. Methode [270]. miracle [37]. Misconception [155]. Miskenning [155]. Misserfolg [156]. modern [36]. moderne [36]. Moeder [197]. Moraal [270]. Morality [270]. Morgen [324, 350, 323, 335].

Moritz [44, 76, 77]. Mother [198, 197]. Müller [101]. Müller-Freienfels [101]. Musical [97]. Musiceeren [97]. Mysterie [14]. Mystery [14]. Myth [28]. Mythe [28].

N [301, 300]. Naamlooze [90]. Naarden [339]. Nameless [90]. Natur [316, 325]. Natural [351]. Nature [316, 325]. Naturwissenschaft [351]. Necessity [275]. Nederlandsch [342]. Neo [81, 321]. Neo-Positivism [321]. Neo-Thomism [81]. Neo-Thomisme [81]. Neurath [172, 192]. Niet [303]. Niet-Leven [303]. Nietzsche [211, 214, 213, 157]. Nietzsches [214]. Nieuwe [224, 25, 5, 159, 98, 50, 36]. Nikolai [151]. Nog [187]. Non [303]. Non-Life [303]. Noodzakelijkheid [275]. North [300]. note [298]. November [260]. Nu [324, 350, 323, 335]. Numbers [95].

Observation [23]. October [342, 259]. old [50, 93]. om [313]. Ondergrond [181]. Onderwijs [178]. Onderzoek [270]. one [65]. Onkenbaarheid [207]. only [65]. ons [313]. Ontmoeten [190]. Ontvangen [236, 273, 327]. Ontwikkeling [180]. Onze [211, 277, 29]. Onzekerheidsrelatie [216]. Onzen [262, 302, 314, 263, 182]. Oorlog [282]. Oorzakelijkheid [46]. Oorzakelijkheidsgeloof [131]. op [339, 213, 286]. Operation [346]. Opmerkingen [154, 228]. opname [313]. opposite [51]. oproep [342]. Optimum [305, 296]. Organiastie [297]. Organisatie [286]. Organisation [353]. Organism [240]. Organisme [240]. Organization [286, 287, 353, 296]. Organizational [297]. Orthodox [360]. Orthodoxer [360]. other [74]. Otto [172, 192]. Oude [93, 50]. Our [211, 277, 262, 302, 314, 263, 182, 29]. overall [266].

path [349]. Pensmaniero [145]. perception [217, 82, 191]. perceptual [191]. Personality [152, 163, 153, 164]. Persönlichkeit [153, 164]. Persoonlijkheid [152, 163]. Perspectieven [32, 8, 17]. Perspectives [32, 8, 17]. Pessimistic [208, 209]. Pessimistische [209]. Pessimistischer [208]. Phenomena [13]. Philosofie [73]. philosopher [337]. Philosophical [315, 316, 325, 230, 231]. Philosophie [117, 188, 300, 143, 58]. Philosophiques [231]. Philosophischen [316, 325]. Philosophy [57, 265, 117, 3, 188, 181, 300, 229, 73, 301, 11, 15, 80, 143, 77, 178, 250, 90, 58]. Phrases [209, 208]. Physica [50, 224, 10, 36, 10]. Physics [224]. Plaats [238]. Place [238]. Plant [24, 39, 53]. Plato [3, 11, 15]. Plotinus [310, 311]. Poesie [293]. Poetry [293]. points [355]. Poljana [85]. Populariseeren [20]. Popularization [20]. Positivism [321]. Positivismus [321]. pour [271]. power [287, 286]. Prickly [24, 39, 53]. Prikkelverschijnselen [24, 39, 53]. Principes [231]. principle [4]. Principles [230, 231]. Priori [129]. Probleem [116, 233, 343, 119, 94, 99, 106, 14]. Probleemstudie [289]. Problem [116, 233, 344, 343, 266, 289, 119, 107, 106, 14, 232, 316, 325, 94, 99]. problemen [10]. Problems [10, 316, 325]. Production [286]. Productiviteit [297]. Productivity [297, 305, 296]. Prof. [249, 339, 298].

Proverbial [2, 9]. Psikologia [145]. Psychoanalytic [184, 206, 234, 317]. Psychoanalytische [184, 206, 234, 317]. Psychological [29, 188, 181, 8, 17, 32, 145]. Psychologie [238, 218, 271, 332, 199, 293]. Psychologique [188]. Psychologisch [32]. Psychologische [8, 17, 29, 181]. Psychology [238, 218, 271, 329, 332, 199, 119, 293].

Questions [262, 302, 284, 314, 333].

Rational [267]. Rationeel [267]. ratios [141]. Rautenstrauch [298]. Reality [169, 226, 137]. received [236, 273, 327]. record [313]. Reflections [162]. regarding [298]. Relatie [224]. Relatie-Leer [224]. Relations [285]. Relationship [224, 216]. Relationships [132, 133, 140]. Religie [108]. Religieus [263]. Religieus-Wijsgeerig [263]. Religieuse [229]. Religieuze [262, 302, 314]. Religion [108]. Religious [262, 302, 314, 263, 229]. Religious-Wise [263]. remarks [191]. requests [313]. Research [270]. Review [26, 43, 89, 193, 201, 244, 251, 279, 272, 171, 210]. Richard [101]. Rijksmuseum [313]. role [296]. Roman [182]. Roomsch [182]. Roomsch-Katholieke [182]. Ruimte [132, 133, 140]. Ruimte-Verhoudingen [132, 133, 140].

Sage [347]. Samenleving [240]. Samenvatting [64]. Sanguinis [127, 121, 126]. Sätze [208]. Schaduwen [7]. Schaffenden [156]. Scheppenden [155]. Schijn [137]. Schlick [44, 76, 77]. Scholen [8, 17, 32]. schools [8, 17, 32]. Schoonheid [116, 30]. Science [253, 254, 255, 256, 257, 259, 260, 261, 320, 354, 357, 358, 362, 190, 174, 175, 321, 108, 266, 276, 351, 173, 183, 361, 359, 322]. sciences [107]. Scripts [179]. Section [288]. Seed [139]. Seele [308]. Sein [338]. Self [217]. Self-perception [217]. Sensation [23]. sensations [278]. Sense [209]. Sensualist [218]. Sensualistische [218]. sentences [191]. September [257]. set [355]. Shadows [7]. shape [141]. Shapes [169]. Shine [137]. side [74]. Signific [290]. Significa [70, 288, 341]. Significance [270, 258, 77, 341, 340]. Significant [116, 212, 177, 185, 339, 269, 328, 227, 160, 109, 125, 134, 289]. signifier [71]. Signifiese [109, 125, 134]. Signifik [340, 71]. Signifika [270]. Signifiker [258]. Signifisch [116]. Signifische [160, 212, 177, 185, 339, 269, 289, 328, 227]. Sin [86]. Sine [127, 121, 126]. Sinn [208]. Sir [35]. Slogans [149]. Social [284]. Sociale [284]. Society [240, 242]. Sociologie [138]. Sociology [138]. Sollsaetze [355]. some [32, 8, 17, 154, 191]. Something [224]. Soul [307, 308]. Space [132, 133, 140]. Space-Relationships [132, 133, 140]. speech [93, 2, 9]. Spel [199]. Spirit [205]. Spiritual [226]. Spranger [238, 239]. Stand [225]. standpoint [107]. Standpunkte [107]. Still [187]. Stof [205]. Stofverschijnselen [168]. Strekking [288]. strijd [80]. String [288]. Structuren [276]. Structures [276]. struggle [80]. Studiegroep [185, 269, 227, 212, 177]. Study [212, 177, 185, 271, 269, 289, 227]. subject [51]. Success [155, 156]. Sudden [282]. summary [64]. Summer [339, 328]. Supernumerary [164]. Swedenborg [166, 167]. symptoms [24, 39, 53]. Synthese [247, 119, 135, 174, 175, 346]. Synthesis [247, 119, 135, 174, 175]. synthetic [93, 18, 22, 34, 41, 48, 54, 61, 68, 88, 105, 170]. Synthetische [18, 22, 34, 41, 48, 54, 61, 68, 88, 105, 170, 93]. System [214].

Taak [288]. Taal [294, 270]. Taalgradatie [228]. Taalkundige [332]. Task [288]. te [339, 313]. Teaching [241]. Techniek [334]. Technique [334]. Teeken [136]. Tegen [124, 80]. tegendeelen [51]. Tegenwoodrige [225]. Teleologie [233, 232]. Teleology [233, 232]. Tendencies [258]. Theoretical [114]. Theoretisch [114]. Theorie [219, 75]. Theory [220, 219, 75, 220]. these [276]. thing [189, 196, 215]. Thinking [263, 176, 315, 213, 16, 267, 145, 74, 292, 21, 138]. Thomism [81]. Thomisme [81]. those [11, 15]. Thought [225, 81]. Thoughts [18, 22, 34, 41, 48, 54, 61, 68, 88, 105, 170, 147, 165]. Three [169]. Tijd [211, 277, 262, 302, 314, 263, 139, 182, 29]. Tijdschriften [243, 252, 274, 280, 291, 299, 306, 312, 318, 326, 336, 345]. Time [211, 277, 262, 302, 314, 263, 139, 182, 29]. Tolstoj [85]. Tolstoy [85]. tomorrow [324, 350, 323, 335]. Top [163]. Tot [233, 218, 199, 135, 91]. Transcendental [315]. Transcendentale [315]. Transit [158]. Trend [225]. Trinity [102]. Truth [28, 66, 82]. Tusschen [130, 97]. Tweeërlei [147]. Two [147]. Type [241]. Typenleer [241].

Überpersönlichen [153, 164]. Ueber [321]. Uit [292, 135, 184, 206, 234, 317, 278]. Uncertainty [216]. Understand [219]. Understanding [29, 220, 66]. Unification [258]. Unity [253, 173, 183, 361, 359, 254, 255, 256, 257, 259, 260, 261, 354, 357, 358, 362, 174, 175]. Universiteit [180]. University [180]. unknown [92].

value [93]. Values [152, 163, 153, 164]. Vereinheitlichungstendenzen [258]. vergelijking [3, 11, 15]. Verhoudingen [132, 133, 140]. Verklaring [168]. Verkrijgen [297]. Verlichting [123]. verscheen [342]. verschil [97]. Verstaan [219]. Verständnistheorie [220]. versus [11, 15]. Verwondering [111]. verzoekt [313]. vierde [141]. violence [149]. Visie [213]. Vision [213]. Vitalisme [6, 87, 112, 148]. Vitality [87, 112, 148, 6]. volgende [313, 342]. Volgt [247]. volkssproken [93]. vom [107]. Voor [285, 276, 29, 81, 77]. Voorlichting [123]. vormverhoudingen [141]. Vragen [262, 302, 284, 314, 333]. Vrijheid [343, 275, 94, 106, 99].

waarde [93]. Waarden [152, 163]. Waarheid [28, 66]. Waarneming [23]. Wahrnehmung [191]. Wahrnehmungssätzen [191]. walk [21]. War [282]. Wat [129, 187, 204, 283, 124]. Wereld [36, 102, 283]. Wereld-beeld [36]. Wereldbeschouwing [333]. Werkelijkheid [169, 137]. Werken [239]. Werte [153, 164]. Wetenschap [276, 190, 108, 320, 322]. Wezen [52, 143, 118, 137]. which [247]. Whitehead [301, 300]. Wijsbegeerte [57, 178, 250, 181, 229, 90, 3, 11, 15, 80, 77]. Wijsgeer [347, 337, 85]. Wijsgeerig [263, 315]. Wijsgeerige [230]. Will [129]. Wirklichkeit [226]. Wisdom [85]. Wise [263]. Wiskunde [116, 72]. Wissenschaft [321]. Wissenschaften [107]. Wonder [37]. Woord [130, 66]. woorden [97]. Word [130, 66]. words [97]. Wordt [129]. Works [239]. World [102, 283, 36]. Worldview [333]. written [236, 273, 327].

Yesterday [7].

Zaak [136]. Zedelijke [282, 139, 186]. Zekerheden [83]. Zelfbeleving [217]. Ziel [307]. Zielkunde [119]. zijde [74]. zijn [51]. Zijnsvormen [169]. Zin [209]. Zinnen [209]. Zinzeggende [2, 9]. Zomerconferentie [339, 328]. Zonde [86]. Zum [232, 191]. Zur [355, 71, 321].

References

Anonymous:1936:IDI

[1] Anonymous. Inleiding. (Dutch) [Introduction]. Synthese, 1(1):1–6, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272768.

Schoenmaekers:1936:ZBDa

[2] Dr. M. H. J. Schoenmaekers. Zinzeggende beeldspraak. (Dutch) [Proverbial image speech]. Synthese, 1(1):7–10, December 1936. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272769.

Labberton:1936:KHWa

[3] Dr. D. van Hinloopen Labberton. De kategorieënleer der Hindhuwijsbegeerte in vergelijking met die van Plato, Aristoteles en Kant. (Dutch) [Category learning captures the Hindu philosophy as compared to that of Plato, Aristotle and Kant]. Synthese, 1(1):11–14, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01272770.

Groot:1936:CVC

[4] Dr. H. Groot. De crisis (?) van het causaliteitsbeginsel. (Dutch) [The crisis (?) of the causality principle]. Synthese, 1(1):15–17, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01272771.

Koch:1936:NED

[5] Dr. H. Koch. Nieuwe economie. (Dutch) [New economy]. Synthese, 1 (1):18–20, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272772.

Kruseman:1936:VMD

[6] Dr. W. M. Kruseman. Vitalisme en mechanisme. (Dutch) [Vitality and mechanism]. Synthese, 1(1):21–24, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272773.

Mannoury:1936:SVG

[7] Prof. G. Mannoury. In de schaduwen van gisteren. (Dutch) [In the shadows of yesterday]. Synthese, 1(1):25–29, December 1936. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272774.

Godefroy:1936:PVEa

[8] Dr. J. C. L. Godefroy. Perspectieven van enkele hedendaagsche psychologische scholen. (Dutch) [Perspectives of some contemporary psychological schools]. Synthese, 1(1):29–33, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272775.

Schoenmaekers:1936:ZBDb

[9] Dr. M. H. J. Schoenmaekers. Zinzeggende beeldspraak. (Dutch) [Proverbial image speech]. Synthese, 1(1):34–39, December 1936. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272776.

Groot:1936:MPP

[10] Dr. H. Groot. Metaphysische problemen der Physica. (German) [Metaphysical problems of physica]. Synthese, 1(1):40–44, December 1936. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272777.

Labberton:1936:KHWb

[11] Dr. D. van Hinloopen Labberton. De kategorieënleer der Hindhuwijsbegeerte in vergelijking met die van Plato, Aristoleles en Kant. (Dutch) [Category learning of the Hindu philosophy versus those of Plato, Aristotle and Kant]. Synthese, 1(1):44–49, December 1936. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272778.

Anonymous:1936:CDCa

[12] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1 (1):49–52, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272779.

Jordan:1936:LLD

[13] Prof. Dr. H. J. Jordan. Leven en levensverschijnselen. (Dutch) [Life and life phenomena]. Synthese, 1(1):53–65, December 1936. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272780.

Schoenmaekers:1936:MPD

[14] Dr. M. H. J. Schoenmaekers. Mysterie en probleem. (Dutch) [Mystery and problem]. Synthese, 1(1):65–70, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272781.

Labberton:1936:KHWc

[15] Dr. D. van Hinloopen Labberton. De kategorieënleer der Hindhuwijsbegeerte in vergelijking met die van Plato, Aristoleles en Kant. (Dutch) [Category learning of the Hindu philosophy versus those of Plato, Aristotle and Kant]. Synthese, 1(1):71–75, December 1936. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272782.

Groot:1936:AVD

[16] Dr. H. Groot. De apologie van het denken. (Dutch) [The apology of thinking]. Synthese, 1(1):75–79, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272783.

Godefroy:1936:PVEb

[17] Dr. J. C. L. Godefroy. Perspectieven van enkele hedendaagsche psychologische scholen. (Dutch) [Perspectives of some contemporary psychological schools]. Synthese, 1(1):79–82, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272784.

Anonymous:1936:SGBa

[18] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):83–88, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272785.

Anonymous:1936:CDCb

[19] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1 (1):89–92, December 1936. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272786.

Schoenmaekers:1936:PCD

[20] Dr. M. H. J. Schoenmaekers. Populariseeren en culturiseeren. (Dutch) [Popularization and culture]. Synthese, 1(1):93–98, December 1936. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272787.

Groot:1936:DGV

[21] Dr. H. Groot. Demoeizame gang van het denken. (Dutch) [Demanding walk of thinking]. Synthese, 1(1):99–106, December 1936. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272788.

Anonymous:1936:SGBb

[22] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):106–112, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272789.

Anonymous:1936:GWB

[23] Anonymous. Gewaarwording, waarneming, beleving. (Dutch) [Sensation, observation, experience]. Synthese, 1(1):113–117, December 1936. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272790.

Kruseman:1936:PBPa

[24] Dr. W. M. Kruseman. Prikkelverschijnselen bij plant, dier en mensch. (Dutch) [Prickly symptoms in plant, animal and man]. Synthese, 1(1): 118–120, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272791.

Koch:1936:NE

[25] Dr. H. Koch. Nieuwe economie. (Dutch) [New economy]. Synthese, 1 (1):121–123, December 1936. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272792.

Anonymous:1936:Ba

[26] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 1(1): 124, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272793.

Anonymous:1936:LC

[27] Anonymous. Leo Chestov. Synthese, 1(1):125–126, December 1936. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272794.

Chestov:1936:MWF

[28] Prof. Dr. Leo Chestov. Mythe en waarheid. (Dutch) [Myth and truth]. Synthese, 1(1):127–135, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01272795.

Mannoury:1936:BVP

[29] Prof. G. Mannoury. De betekenis van de psychologische begripsontleding voor onze tijd. (Dutch) [The meaning of the psychological understanding analysis for our time]. Synthese, 1(1):136–138, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272796.

Schoenmaekers:1936:S

[30] Dr. M. H. J. Schoenmaekers. Schoonheid. (Dutch) [Beauty]. Synthese, 1 (1):139–142, December 1936. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272797.

Groot:1936:BVMa

[31] Dr. H. Groot. De bouw van de materie. (Dutch) [The construction of the matter]. Synthese, 1(1):143–147, December 1936. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272798.

Godefroy:1936:PVE

[32] Dr. J. C. L. Godefroy. Perspectieven van enkele hedendaagsche Psychologisch Scholen. (Dutch) [Perspectives of some contemporary Psychological Schools]. Synthese, 1(1):148–151, December 1936. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272799.

Anonymous:1936:Ca

[33] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1(1): 151, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272800.

Anonymous:1936:SGBc

[34] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):152–156, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272801.

Anonymous:1936:SJH

[35] Anonymous. Sir James H. Jeans. Synthese, 1(1):157–159, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272802.

Jeans:1936:NWB

[36] Sir James H. Jeans. Het nieuwe wereld-beeld van de moderne physica. (Dutch) [The new world image of modern physica]. Synthese, 1(1):160– 169, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272803.

Schoenmaekers:1936:W

[37] Dr. M. H. J. Schoenmaekers. Het wonder. (Dutch) [The miracle]. Synthese, 1(1):170–174, December 1936. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272804.

Groot:1936:BVMb

[38] Dr. H. Groot. De bouw van de materie. (Dutch) [The construction of matter]. Synthese, 1(1):175–179, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272805.

Kruseman:1936:PBPb

[39] Dr. W. M. Kruseman. Prikkelverschijnselen bij plant, dier en mensch. (Dutch) [Prickly symptoms in plant, animal and man]. Synthese, 1(1): 180–183, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272806.

Anonymous:1936:Ea

[40] Anonymous. Erratum. Synthese, 1(1):183, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272807;   http://link.
springer.com/content/pdf/10.1007/BF01272807.pdf.

Anonymous:1936:SGBd

[41] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):184–189, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272808.

Anonymous:1936:Cb

[42] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1(1): 190, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272809.

Anonymous:1936:Bb

[43] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 1(1):191– 192, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272810.

Anonymous:1936:MSa

[44] Anonymous. Moritz Schlick. Synthese, 1(1):193–201, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272811.

Anonymous:1936:EHG

[45] Anonymous. Eduard von Hartmann. Synthese, 1(1):202–206, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01272812.

Schoenmaekers:1936:O

[46] Dr. M. H. J. Schoenmaekers. Oorzakelijkheid. (Dutch) [Causality]. Synthese, 1(1):207–214, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01272813.

Groot:1936:BVMc

[47] Dr. H. Groot. De bouw van de materie. (Dutch) [The construction of matter]. Synthese, 1(1):215–221, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272814.

Anonymous:1936:SGBe

[48] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):221–226, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272815.

Anonymous:1936:Cc

[49] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1 (1):227–228, December 1936. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272816.

Anonymous:1936:ONP

[50] Anonymous. De oude en de nieuwe physica. (Dutch) [The old and the new physica]. Synthese, 1(1):229, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272817.

Schoenmaekers:1936:SZT

[51] Dr. M. H. J. Schoenmaekers. Het subject en zijn tegendeelen. (Dutch) [The subject and its opposite]. Synthese, 1(1):230–246, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272818.

Groot:1936:WVM

[52] Dr. H. Groot. Het wezen van de materie. (Dutch) [The essence of matter]. Synthese, 1(1):247–251, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01272819.

Kruseman:1936:PBPc

[53] Dr. W. M. Kruseman. Prikkelverschijnselen bij plant, dier en mensch. (Dutch) [Prickly symptoms in plant, animal and man]. Synthese, 1(1): 252–255, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272820.

Anonymous:1936:SGBf

[54] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):256–258, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272821.

Anonymous:1936:Cd

[55] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1(1): 259, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272822.

Anonymous:1936:Eb

[56] Anonymous. Errata. Synthese, 1(1):259, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272823;   http://link.
springer.com/content/pdf/10.1007/BF01272823.pdf.

Anonymous:1936:HW

[57] Anonymous. Hedendaagsche wijsbegeerte. (Dutch) [Contemporary philosophy]. Synthese, 1(1):261, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272824.

Wolf:1936:PCF

[58] Dr. Herman Wolf. Philosophie en cultuur. (Dutch) [Philosophy and culture]. Synthese, 1(1):262–274, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01272825.

Elias:1936:KLE

[59] J. S. Witsen Elias. De kunstgeschiedenis in het licht eener algemeene geesteswetenschap. (Dutch) [The history of art in the light of general humanities]. Synthese, 1(1):275–281, December 1936. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272826.

Schoenmaekers:1936:G

[60] Dr. M. H. J. Schoenmaekers. Het gebaar. (Dutch) [The gesture]. Synthese, 1(1):282–287, December 1936. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272827.

Anonymous:1936:SGBg

[61] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):288–292, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272828.

Anonymous:1936:Ce

[62] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1(1): 292, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272829.

Anonymous:1936:MB

[63] Anonymous. Martin Buber. Synthese, 1(1):293–299, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272830.

Anonymous:1936:KS

[64] Anonymous. Korte samenvatting. (Dutch) [Brief summary]. Synthese, 1 (1):299–300, December 1936. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272831.

Buber:1936:EEG

[65] Prof. Dr. Martin Buber. “Der Einzige” und “der Einzelne”. (German) [“The only one” and “the individual”]. Synthese, 1(1):300–308, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01272832.

Schoenmaekers:1936:WBW

[66] Dr. M. H. J. Schoenmaekers. Woord, begrip, waarheid. (Dutch) [Word, understanding, truth]. Synthese, 1(1):309–313, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272833.

Hartzfeld:1936:IM

[67] Mr. C. A. J. Hartzfeld. Intuïtieve menschenkennis. (Dutch) [Intuitive knowledge]. Synthese, 1(1):313–318, December 1936. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01272834.

Anonymous:1936:SGBh

[68] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 1(1):319–322, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272835.

Anonymous:1936:Cf

[69] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1 (1):323–324, December 1936. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01272836.

Anonymous:1936:S

[70] Anonymous. Significa. Synthese, 1(1):325–339, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272837.

Mannoury:1936:BSG

[71] Prof. G. Mannoury. Zur Begriffsbestimmung der Signifik. (German) [The definition of the signifier]. Synthese, 1(1):340–341, December 1936. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272838.

Groot:1936:MGW

[72] Dr. H. Groot. De metaphysische grond der wiskunde. (Dutch) [The metaphysical ground of mathematics]. Synthese, 1(1):342–345, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272839.

deHartog:1936:EPM

[73] Prof. Dr. A. H. de Hartog. Empirie, philosofie, metaphysica. (Dutch) [Empire, philosophy, metaphysics]. Synthese, 1(1):346–350, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272840.

Schoenmaekers:1936:AGZ

[74] Dr. M. H. J. Schoenmaekers. Aan gene zijde van het denken. (Dutch) [On the other side of the thinking]. Synthese, 1(1):351–354, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272841.

Kruseman:1936:GTV

[75] Dr. W. M. Kruseman. De “gestalt”-theorie van Köhler. (Dutch) [The “Gestalt” theory of Köhler]. Synthese, 1(1):355–360, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01272842.

Anonymous:1936:MSb

[76] Anonymous. Moritz Schlick. Synthese, 1(1):361, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272843.

Waismann:1936:BVM

[77] Friedrich Waismann. De beteekenis van Moritz Schlick voor de wijsbegeerte. (Dutch) [The significance of Moritz Schlick for philosophy]. Synthese, 1(1):361–370, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01272844.

Anonymous:1936:Ec

[78] Anonymous. Errata. Synthese, 1(1):370, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01272845;   http://link.
springer.com/content/pdf/10.1007/BF01272845.pdf.

Anonymous:1936:Cg

[79] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 1(1): 370, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272846.

Oldewelt:1936:SWT

[80] Dr. H. M. J. Oldewelt. De strijd der wijsbegeerte tegen de argeloosheid. (Dutch) [The struggle of philosophy against impotence]. Synthese, 1(1): 371–376, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272847.

Robbers:1936:BVN

[81] Dr. H. Robbers. De beteekenis van het neo-Thomisme voor het hedendaagsche Katholieke denkleven. (Dutch) [The meaning of neo-Thomism for contemporary Catholic thought life]. Synthese, 1(1):377– 382, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272848.

Schoenmaekers:1936:CBF

[82] Dr. M. H. J. Schoenmaekers. Constateerwaarheid en belevingswaarheid. (Dutch) [Constant truth and perception of truth]. Synthese, 1(1):383– 390, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01272849.

Groot:1936:CZG

[83] Dr. H. Groot. De crisis der zekerheden. (Dutch) [The crisis of confidences]. Synthese, 1(1):391–400, December 1936. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01272850.

Anonymous:1937:LC

[84] Anonymous. Leo chestov. Synthese, 2(1):2–10, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880390.

Chestov:1937:JPA

[85] Prof. Dr. Leo Chestov. Jasnaja-Poljana en Astapowo Tolstoj als Wijsgeer. (Dutch) [Jasnaja-Poljana and Astapowo Tolstoy as wisdom]. Synthese, 2 (1):11–19, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880391.

Schoenmaekers:1937:ZDS

[86] Dr. M. H. J. Schoenmaekers. Zonde. (Dutch) [Sin]. Synthese, 2(1):20– 23, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880392.

Kruseman:1937:HVVa

[87] Dr. W. M. Kruseman. De historie van het vitalisme. (Dutch) [The history of vitality]. Synthese, 2(1):24–27, December 1937. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880393.

Anonymous:1937:SGBa

[88] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 2(1):28–38, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880394.

Anonymous:1937:BDB

[89] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 2(1):39– 40, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880395.

Schoenmaekers:1937:WND

[90] Dr. M. H. J. Schoenmaekers. Wijsbegeerte, de naamlooze. (Dutch) [Philosophy, the nameless]. Synthese, 2(1):41–45, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880396.

Schachter:1937:BTA

[91] Dr. Josef Schächter. Bijdrage tot de analyse van het begrip “cultuur”. (Dutch) [Contribution to the analysis of the concept “culture”]. Synthese, 2(1):47–54, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880397.

Groot:1937:U

[92] Dr. H. Groot. [unknown]. Synthese, 2(1):55–67, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880398.

Beishuizen:1937:OVH

[93] Thole Beishuizen. Oude volkssproken en haar synthetische waarde. (Dutch) [Old folk’s speech and its synthetic value]. Synthese, 2(1):68– 72, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880399.

Jordan:1937:PVDa

[94] Prof. Dr. H. J. Jordan. Het probleem der vrijheid. (Dutch) [The Freedom Problem]. Synthese, 2(1):73–81, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880400.

Schoenmaekers:1937:HGD

[95] Dr. M. H. J. Schoenmaekers. “Heilige” getallen. (Dutch) [“Holy” numbers]. Synthese, 2(1):82–94, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880401.

Groot:1937:BCD

[96] Dr. H. Groot. Beschouwingen over cosmogonie. (Dutch) [Considerations about cosmogony]. Synthese, 2(1):95–100, December 1937. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880402.

Esch:1937:VTM

[97] W. Esch. Over het verschil tusschen “musiceeren” met woorden en het “bouwen” met woorden. (Dutch) [On the difference between “musical” with words and ‘construction’ with words]. Synthese, 2(1):101– 105, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880403.

Koch:1937:NED

[98] Dr. H. Koch. Nieuwe economie. (Dutch) [New economy]. Synthese, 2(1): 105–108, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880404.

Jordan:1937:PVDb

[99] Prof. Dr. H. J. Jordan. Het Probleem der Vrijheid. (Dutch) [The Freedom Problem]. Synthese, 2(1):109–116, December 1937. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880405.

Anonymous:1937:EDE

[100] Anonymous. Erratum. Synthese, 2(1):116, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880406;   http://link.
springer.com/content/pdf/10.1007/BF00880406.pdf.

Anonymous:1937:RMF

[101] Anonymous. Richard Müller-Freienfels. Synthese, 2(1):117–124, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880407.

Muller-Freienfels:1937:DWD

[102] Prof. Richard Müller-Freienfels. De drieëenheid der wereld. (Dutch) [The trinity of the world]. Synthese, 2(1):125–133, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880408.

Schoenmaekers:1937:BDE

[103] Dr. M. H. J. Schoenmaekers. Beleving. (Dutch) [Experience]. Synthese, 2 (1):134–139, December 1937. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880409.

Groot:1937:BCa

[104] Dr. H. Groot. Beschouwingen over cosmogonie. (Dutch) [Considerations about cosmogony]. Synthese, 2(1):140–145, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880410.

Anonymous:1937:SGBb

[105] Anonymous. De synthetische gedachte in het buitenland. (Dutch) [Synthetic thoughts in abroad]. Synthese, 2(1):146–148, December 1937. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00880411.

Jordan:1937:PVG

[106] Prof. Dr. H. J. Jordan. Het probleem der vrijheid. (Dutch) [The problem of freedom]. Synthese, 2(1):149–154, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880412.

Jordan:1937:PFS

[107] Prof. Dr. H. J. Jordan. Das Problem der Freiheit vom Standpunkte der Empirischen Wissenschaften. (German) [The problem of freedom from the standpoint of empirical sciences]. Synthese, 2(1):155–158, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00880413.

Schachter:1937:RWF

[108] Dr. Josef Schächter. Religie en wetenschap. (Dutch) [Religion and science]. Synthese, 2(1):159–167, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00880414.

Brouwer:1937:SDa

[109] L. E. J. Brouwer, Fred Van Eeden, J. Van Ginneken, and S. J. G. Mannoury. Signifiese dialogen. (Dutch) [Significant dialogue]. Synthese, 2 (1):168–174, December 1937. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880415.

Groot:1937:BCb

[110] Dr. H. Groot. Beschouwingen over cosmogonie. (Dutch) [Considerations about cosmogony]. Synthese, 2(1):175–181, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880416.

Schoenmaekers:1937:VCA

[111] Dr. M. H. J. Schoenmaekers. Verwondering contra angst. (Dutch) [Astonishing contrast anxiety]. Synthese, 2(1):182–184, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880417.

Kruseman:1937:HVVb

[112] Dr. W. M. Kruseman. De historie van het vitalisme. (Dutch) [The history of vitality]. Synthese, 2(1):185–188, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880418.

Anonymous:1937:Ea

[113] Anonymous. Errata. Synthese, 2(1):188, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880419;   http://link.
springer.com/content/pdf/10.1007/BF00880419.pdf.

Jordan:1937:HDT

[114] Prof. Dr. H. J. Jordan. Hans Driesch als theoretisch bioloog. (Dutch) []. (Dutch) [Hans Driesch as theoretical biologist]. Synthese, 2(1):190– 191, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880420.

Driesch:1937:AEa

[115] Prof. Dr. Hans Driesch. Das Apriori als Erlebnis. (German) [The apriori as an experience]. Synthese, 2(1):192–196, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880421.

Anonymous:1937:SWS

[116] Anonymous. De schoonheid der wiskunde als signifisch probleem. (Dutch) [The beauty of mathematics as a significant problem]. Synthese, 2(1):197– 201, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880422.

Jaspers:1937:EP

[117] Karl Jaspers. De existentie-philosophie. (Dutch) [The existence philosophy]. Synthese, 2(1):202–209, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880423.

Schoenmaekers:1937:WVK

[118] Dr. M. H. J. Schoenmaekers. Over het wezen van het Katholicisme. (Dutch) [On the essence of Catholicism]. Synthese, 2(1):210– 222, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880424.

Hartzfeld:1937:PVS

[119] Mr. C. A. J. Hartzfeld. Over een probleem van synthese in de zielkunde. (Dutch) [About a problem of synthesis in psychology]. Synthese, 2(1): 223–228, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880425.

Anonymous:1937:Eb

[120] Anonymous. Erratum. Synthese, 2(1):228, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880426;   http://link.
springer.com/content/pdf/10.1007/BF00880426.pdf.

Chestov:1937:SESa

[121] Prof. Dr. Leo Chestov. Sine effusione sanguinis. Synthese, 2(1):230– 244, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880427.

Driesch:1937:AEb

[122] Prof. Dr. Hans Driesch. Das apriori als erlebnis. (German) [The apriori as an experience]. Synthese, 2(1):246–254, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880428.

Schoenmaekers:1937:VIV

[123] Dr. M. H. J. Schoenmaekers. Voorlichting, inlichting, verlichting. (Dutch) [Information, information, lighting]. Synthese, 2(1):255–257, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880429.

Esch:1937:WT

[124] W. Esch. Wàt is er tegen?. (Dutch) [What is it against?]. Synthese, 2 (1):258–260, December 1937. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880430.

Brouwer:1937:SDb

[125] L. E. J. Brouwer, Fred van Eeden, J. van Ginneken, and S. J. G. Mannoury. Signifiese dialogen. (Dutch) [Significant dialogue]. Synthese, 2(1):261– 268, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880431.

Chestov:1937:SESb

[126] Prof. Dr. Leo Chestov. Sine effusione sanguinis. Synthese, 2(1):269– 282, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880432.

Anonymous:1937:SES

[127] Anonymous. Sine effusione sanguinis. Synthese, 2(1):283–284, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880433.

Driesch:1937:AEc

[128] Prof. Dr. Hans Driesch. Das Apriori als Erlebnis. (German) [The apriori as an experience]. Synthese, 2(1):285–296, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880434.

Anonymous:1937:WPB

[129] Anonymous. Wat a priori beleefd wordt. (Dutch) [What a priori will be]. Synthese, 2(1):297–298, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880435.

Groot:1937:WBT

[130] Dr. H. Groot. Het woord als brug tusschen mensch en mensch. (Dutch) [The word as bridge between man and man]. Synthese, 2(1):299– 308, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880436.

Schoenmaekers:1937:EBO

[131] Dr. M. H. J. Schoenmaekers. De emotie bij het oorzakelijkheidsgeloof. (Dutch) [The emotion at the causality faith]. Synthese, 2(1):309–310, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880437.

Anonymous:1937:BRV

[132] Anonymous. Beschouwingen over ruimte-verhoudingen. (Dutch) [Considerations on space-relationships]. Synthese, 2(1):311–312, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00880438.

Jodjana:1937:BRVa

[133] Raden Ajoe Jodjana. Beschouwingen over ruimte-verhoudingen. (Dutch) [Considerations on space-relationships]. Synthese, 2(1):313–315, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880439.

Brouwer:1937:SDc

[134] L. E. J. Brouwer, Fred van Eeden, J. van Ginneken, and S. J. G. Mannoury. Signifiese dialogen. (Dutch) [Significant dialogue]. Synthese, 2(1):316– 324, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880440.

Holstijn:1937:UAT

[135] Dr. A. J. Westerman Holstijn. Uit de antithesen tot de synthese. (Dutch) [From the antithesis to the synthesis]. Synthese, 2(1):325–335, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880441.

Jordan:1937:ZTT

[136] Prof. Dr. Leo Jordan. Zaak en teeken — teeken en zaak. (Dutch) [Case and draw — draw and case]. Synthese, 2(1):336–346, December 1937.

CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00880442.

Wolf:1937:DWS

[137] Dr. Herman Wolf. Droom en werkelijkheid, schijn en wezen. (Dutch) [Dream and reality, shine and being]. Synthese, 2(1):347–363, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880443.

Thomassen:1937:SVD

[138] J. Thomassen. Sociologie van het denken?. (Dutch) [Sociology of thinking?]. Synthese, 2(1):364–370, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880444.

Daan:1937:ZBV

[139] Albert Daan. De zedelijke behoefte van dezen tijd. (Dutch) [The seed of this time]. Synthese, 2(1):371–380, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880445.

Jodjana:1937:BRVb

[140] Raden Ajoe Jodjana. Beschouwingen over ruimte-verhoudingen. (Dutch) [Considerations on space-relationships]. Synthese, 2(1):381–385, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880446.

Anonymous:1937:IMV

[141] Anonymous. III Het meten der vormverhoudingen. De vierde dimensie. (Dutch) [III The measure of the shape ratios. The fourth dimension]. Synthese, 2(1):386–392, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00880447.

Anonymous:1937:Ia

[142] Anonymous. Inleiding. (Dutch) [Introduction]. Synthese, 2(1):393– 394, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880448.

Schachter:1937:WPG

[143] Dr. Josef Schächter. Over het wezen der philosophie. (Dutch) [About the essence of philosophy]. Synthese, 2(1):395–405, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880449.

Anonymous:1937:Ib

[144] Anonymous. Inleiding. (Dutch) [Introduction]. Synthese, 2(1):406– 407, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880450.

Mannoury:1937:PAM

[145] G. Mannoury. Psikologia analizo de la matematika pensmaniero. (Esperanto) [Psychological analysis of mathematical thinking]. Synthese, 2 (1):407–418, December 1937. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880451.

Schoenmaekers:1937:B

[146] Dr. M. H. J. Schoenmaekers. Over bewustzijn. (Dutch) [On consciousness]. Synthese, 2(1):419–428, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00880452.

Hartzfeld:1937:TD

[147] Mr. C. A. J. Hartzfeld. Over tweeërlei denken. (Dutch) [About two thoughts]. Synthese, 2(1):429–437, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880453.

Kruseman:1937:HVVc

[148] Dr. W. M. Kruseman. De historie van het vitalisme. (Dutch) [The history of vitality]. Synthese, 2(1):438–442, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880454.

Esch:1937:LG

[149] W. Esch. Leuzen als geweldmiddel. (Dutch) [Slogans as means of violence]. Synthese, 2(1):443–448, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00880455.

Anonymous:1937:Ec

[150] Anonymous. Erratum. Synthese, 2(1):448, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880456;   http://link.
springer.com/content/pdf/10.1007/BF00880456.pdf.

Anonymous:1937:NB

[151] Anonymous. Nikolai Berdjajew. Synthese, 2(1):450–455, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00880457.

Berdjajew:1937:MPBa

[152] Prof. Nikolai Berdjajew. De menschelijke persoonlijkheid en de bovenpersoonlijke waarden. (Dutch) [Human personality and extraordinary values]. Synthese, 2(1):456–464, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880458.

Berdjajew:1937:MPUa

[153] Prof. Nikolai Berdjajew. Die Menschliche Persönlichkeit und Die Überpersönlichen Werte. (German) [Human personality and extraordinary values]. Synthese, 2(1):465–466, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880459.

Anonymous:1937:EOG

[154] Anonymous. Eenige opmerkingen over het genie. (Dutch) [Some comments about the genius]. Synthese, 2(1):467–470, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880460.

Hirschhorn:1937:SMV

[155] Dr. S. Hirschhorn. Over success en miskenning van den scheppenden mensch. (Dutch) [About success and misconception of man the creator]. Synthese, 2(1):471–476, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00880461.

Hirschhorn:1937:EMS

[156] Dr. S. Hirschhorn. Über Erfolg und Misserfolg des Schaffenden. (German) [Success and Failure of the Creator]. Synthese, 2(1):477–478, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880462.

Rispens:1937:FN

[157] J. A. Rispens. Over Friedrich Nietzsche. (Dutch) [About Friedrich Nietzsche]. Synthese, 2(1):479–485, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880463.

Groot:1937:MLD

[158] Dr. H. Groot. Metafysika en logistiek als doorgangsstadia. (Dutch) [Metaphysics and logistics as transit cities]. Synthese, 2(1):486–491, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880464.

Koch:1937:NE

[159] Dr. H. Koch. Nieuwe economie. (Dutch) [New economy]. Synthese, 2 (1):492–494, December 1937. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880465.

Anonymous:1937:SK

[160] Anonymous. Signifische klacht. (Dutch) [Significant complaint]. Synthese, 2(1):495, December 1937. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00880466.

Anonymous:1937:Ed

[161] Anonymous. Errata. Synthese, 2(1):496, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880467;   http://link.
springer.com/content/pdf/10.1007/BF00880467.pdf.

Anonymous:1937:BK

[162] Anonymous. Beschouwingen over kunstgeschiedenis. (Dutch) [Reflections on art history]. Synthese, 2(1):497–502, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880468.

Berdjajew:1937:MPBb

[163] Prof. Nikolai Berdjajew. De menschelijke persoonlijkheid en de bovenpersoonlijke waarden. (Dutch) [The human personality and the top values]. Synthese, 2(1):503–509, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00880469.

Berdjajew:1937:MPUb

[164] Prof. Nikolai Berdjajew. Die Menschliche Persönlichkeit und Die Überpersönlichen Werte. (German) [The Human Personality and The Supernumerary Values]. Synthese, 2(1):510–511, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880470.

Schoenmaekers:1937:EG

[165] Dr. M. H. J. Schoenmaekers. Eenzame gedachten. (Dutch) [Lonely thoughts]. Synthese, 2(1):512–513, December 1937. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00880471.

vanOs:1937:SKa

[166] Prof. Dr. C. H. van Os. Swedenborg and Kant. Synthese, 2(1):514– 525, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880472.

vanOs:1937:SKb

[167] Prof. Dr. C. H. van Os. Swedenborg and Kant. Synthese, 2(1):525– 526, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880473.

Tombrock:1937:VVS

[168] Dr. W. Tombrock. Over de verklaring van stofverschijnselen. (Dutch) [About the declaration of diseases]. Synthese, 2(1):527–528, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880474.

Dietz:1937:DZE

[169] Dr. P. A. Dietz. De drie zijnsvormen der empirische werkelijkheid. (Dutch) [The three shapes of empirical reality]. Synthese, 2(1):529–538, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00880475.

Anonymous:1937:SGBc

[170] Anonymous. De synthetische gedachten in het buitenland. (Dutch) [Synthetic thoughts abroad]. Synthese, 2(1):539–541, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00880476.

Mannoury:1937:B

[171] G. Mannoury. Boekbespreking. (Dutch) [Book review]. Synthese, 2(1): 542–543, December 1937. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00880477.

Anonymous:1938:ONa

[172] Anonymous. Otto Neurath. Synthese, 3(1):1–9, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930205.

Anonymous:1938:IIUa

[173] Anonymous. International Institute for the Unity of Science. Synthese, 3 (1):9, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930206.

Neurath:1938:EESa

[174] Dr. Otto Neurath. Eenheidswetenschap als empiristische synthese. (Dutch) [Unity science as empirical synthesis]. Synthese, 3(1):10–17, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930207.

Neurath:1938:EESb

[175] Otto Neurath. Einheitswissenschaft Als Empiristische Synthese. (German) [Unity science as empirical synthesis]. Synthese, 3(1):18–19, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00930208.

Buys:1938:DDA

[176] Dr. W. R. van Brakell Buys. Over het denken. (Dutch) [About thinking]. Synthese, 3(1):20–30, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930209.

Anonymous:1938:ISSa

[177] Anonymous. Internationale Signifische Studiegroep. (Dutch) [International Significant Study Group]. Synthese, 3(1):30, December 1938. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930210.

Anonymous:1938:WOD

[178] Anonymous. Wijsbegeerte en onderwijs. (Dutch) [Philosophy and education]. Synthese, 3(1):31–33, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00930211.

Anonymous:1938:IBG

[179] Anonymous. Ingekomen boeken en geschriften. (Dutch) [Books and scripts arrived]. Synthese, 3(1):33, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00930212.

Clay:1938:OVU

[180] Prof. Dr. J. Clay. De ontwikkeling van de universiteit. (Dutch) [The development of the university]. Synthese, 3(1):34–48, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930213.

Oldewelt:1938:POV

[181] Dr. H. M. J. Oldewelt. De psychologische ondergrond Van Bergson’s wijsbegeerte. (Dutch) [The psychological background of Bergson’s philosophy]. Synthese, 3(1):49–61, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00930214.

Schoenmaekers:1938:RKA

[182] Dr. M. H. J. Schoenmaekers. Roomsch-Katholieke apologetica in onzen tijd. (Dutch) [Roman Catholic apologetics in our time]. Synthese, 3(1): 62–68, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930215.

Anonymous:1938:IIUb

[183] Anonymous. International Institute for the Unity of Science. Synthese, 3(1):68, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930216.

Holstijn:1938:HUPa

[184] Dr. A. J. Westerman Holstijn. Hoofdstukken uit de psychoanalytische criminologie. (Dutch) [Chapters from psychoanalytic criminology].

Synthese, 3(1):69–81, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930217.

Anonymous:1938:ISSb

[185] Anonymous. Internationale signifische studiegroep. (Dutch) [International significant study group]. Synthese, 3(1):81, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930218.

Daan:1938:ZDD

[186] Albert Daan. De zedelijke dynamiek. (Dutch) [The divine dynamics]. Synthese, 3(1):82–92, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930219.

Anonymous:1938:MNW

[187] Anonymous. Over metaphysica en nog wat. (Dutch) [About metaphysics and still what]. Synthese, 3(1):93–98, December 1938. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930220.

Oldewelt:1938:FPP

[188] Dr. H. M. J. Oldewelt. Le fond psychologique de la philosophie bergsonienne. (French) [The psychological background of Bergsonian philosophy]. Synthese, 3(1):99–101, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930221.

Driesch:1938:Da

[189] Prof. Dr. Hans Driesch. I. Das “Ding”. (German) [I. The “thing”]. Synthese, 3(1):102–105, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930222.

Kappers:1938:OVG

[190] C. U. Ariëns Kappers. Het ontmoeten van geloof en wetenschap. (Dutch) [Meetings of faith and science]. Synthese, 3(1):106–117, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930223.

Steinhardt:1938:EBB

[191] Dr. Käthe Steinhardt. Einige Bemerkungen zum Begriffe der Wahrnehmung und zu den Wahrnehmungssätzen. (German) [Some remarks on the concept of perception and perceptual sentences]. Synthese, 3(1):118–127, December 1938. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930224.

Anonymous:1938:ONb

[192] Anonymous. Otto Neurath. Synthese, 3(1):128, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930225.

Anonymous:1938:Ba

[193] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 3(1):129– 131, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930226.

Anonymous:1938:Ma

[194] Anonymous. Mededeelingen. (Dutch) [Mediations]. Synthese, 3(1): 132, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930227.

Chestov:1938:AJ

[195] Prof. Dr. Leo Chestov. Athene en jeruzalem. (Dutch) [Athens and Jerusalem]. Synthese, 3(1):133–135, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930228.

Driesch:1938:Db

[196] Prof. Dr. Hans Driesch. Das “ding”. (German) [The “thing”]. Synthese, 3 (1):136–143, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930229.

Buys:1938:CVK

[197] Dr. W. R. van Brakell Buys. De conceptie van Kali de moeder. (Dutch) [The conception of Kali the mother]. Synthese, 3(1):144–157, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00930230.

Buys:1938:CKM

[198] Dr. W. R. van Brakell Buys. Le culte de Kali, la mère. (French) [The cult of Kali, the mother]. Synthese, 3(1):158, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930231.

Capp:1938:ITP

[199] Prof. Dr. E. A. D. E. Capp. Inleiding tot de psychologie van het spel. (Dutch) [Introduction to the psychology of the game]. Synthese, 3(1): 159–172, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930232.

Groot:1938:C

[200] Dr. H. Groot. Cosmogonie. (Dutch) [Cosmogonie]. Synthese, 3(1):173– 187, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930233.

Anonymous:1938:Bb

[201] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 3(1):188– 191, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930234.

Anonymous:1938:C

[202] Anonymous. Correspondentie. (Dutch) [Correspondence]. Synthese, 3(1): 192, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930235.

Anonymous:1938:Mb

[203] Anonymous. Mededeelingen. (Dutch) [Mediations]. Synthese, 3(1): 192, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930236.

Schoenmaekers:1938:WBG

[204] Dr. M. H. J. Schoenmaekers. Wat beteekent “gelooven”?. (Dutch) [What does “believe” mean?]. Synthese, 3(1):193–196, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930237.

vanLohuizen:1938:GS

[205] Dr. T. van Lohuizen. Geest en stof. (Dutch) [Spirit and dust]. Synthese, 3(1):197–209, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930238.

Holstijn:1938:HUPb

[206] Dr. A. J. Westerman Holstijn. Hoofdstukken uit de psychoanalytische criminologie. (Dutch) [Chapters from psychoanalytic criminology]. Synthese, 3(1):210–218, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930239.

Anonymous:1938:KO

[207] Anonymous. Over “kenbaarheid” en “onkenbaarheid’. (Dutch) [About “knowability” and “indifference ’]. Synthese, 3(1):219–222, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00930240.

Schachter:1938:SPS

[208] Dr. Josef Schächter. Der Sinn Pessimistischer Sätze. (German) [The Meaning of Pessimistic Phrases]. Synthese, 3(1):223–233, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930241.

Schachter:1938:ZVP

[209] Dr. Josef Schächter. De “zin” van pessimistische zinnen. (Dutch) [The “sense” of pessimistic phrases]. Synthese, 3(1):234, December 1938. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930242.

Mannoury:1938:B

[210] G. Mannoury. Boekbespreking. (Dutch) [Book review]. Synthese, 3(1): 235–236, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930243.

Anonymous:1938:FNO

[211] Anonymous. Friedrich Nietzsche en onze tijd. (Dutch) [Friedrich Nietzsche and our time]. Synthese, 3(1):237–257, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930244.

Anonymous:1938:ISS

[212] Anonymous. Internationale Signifische Studiegroep. (Dutch) [International Significant Study Group]. Synthese, 3(1):257, December 1938.

CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930245.

Drescher:1938:VVK

[213] Dr. Wilhelmine Drescher. De visie van Kurt Hildebrandt op Nietzsche’s denken. (Dutch) [The vision of Kurt Hildebrandt on Nietzsche’s thinking]. Synthese, 3(1):259–270, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930246.

Drescher:1938:KHD

[214] Dr. Wilhelmine Drescher. Kurt Hildebrandts Deutung Von Nietzsches “System”. (German) [Kurt Hildebrandt’s interpretation of Nietzsche’s “system”]. Synthese, 3(1):271–275, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930247.

Driesch:1938:Dc

[215] Prof. Dr. Hans Driesch. Das “ding”. (German) [The “thing”]. Synthese, 3 (1):276–279, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930248.

Kruseman:1938:OB

[216] Dr. W. M. Kruseman. De onzekerheidsrelatie in de biologie. (Dutch) [The uncertainty relationship in biology]. Synthese, 3(1):280–285, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00930249.

Janssen:1938:Z

[217] Ir. G. E. M. Janssen. Zelfbeleving. (Dutch) [Self-perception]. Synthese, 3 (1):286–292, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930250.

Anonymous:1938:VSP

[218] Anonymous. Van sensualistische psychologie tot levenspsychologie. (Dutch) [From sensualist psychology to life psychology]. Synthese, 3(1): 293–300, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930251.

Muller-Freienfels:1938:KTV

[219] Prof. Dr. Richard Müller-Freienfels. Kennis-theorie en “verstaan”-theorie. (Dutch) [Knowledge theory and “understand”-theory]. Synthese, 3(1): 301–313, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930252.

Muller-Freienfels:1938:EV

[220] Prof. Dr. Richard Müller-Freienfels. Erkenntnistheorie und Verständnistheorie. (German) [Cognitive Theory and Understanding Theory]. Synthese, 3(1): 314–318, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930253.

Os:1938:CL

[221] Prof. Dr. C. H. Van Os. De crisis der logica. (Dutch) [The crisis der logica]. Synthese, 3(1):319–334, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930254.

Anonymous:1938:CL

[222] Anonymous. The crisis of logic. Synthese, 3(1):335–336, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930255.

Anonymous:1938:Ea

[223] Anonymous. Errata. Synthese, 3(1):336, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930256;   http://link.
springer.com/content/pdf/10.1007/BF00930256.pdf.

Beishuizen:1938:IOR

[224] Thole Beishuizen. Iets over de relatie-leer der Chineezen en de nieuwe physica. (Dutch) [Something about the relationship doctrine of the Chinese and the new physics]. Synthese, 3(1):337–340, December 1938. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930257.

Fahrenfort:1938:TSV

[225] Dr. J. J. Fahrenfort. De tegenwoodrige stand van de evolutie-gedachte in de ethnologie. (Dutch) [The contemporary trend of evolution-thought in ethnology]. Synthese, 3(1):341–355, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930258.

Martin:1938:GFG

[226] Prof. Buber Martin. Die Geistige Forderung und die Geschichtliche Wirklichkeit. (German) [Spiritual demand and historical reality]. Synthese, 3 (1):356–368, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930259.

Mannoury:1938:ISS

[227] G. Mannoury. De Internationale Signifische Studiegroep (I.S.S.). (Dutch) [International Significant Study Group (I.S.S.)]. Synthese, 3(1):369– 371, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930260.

Mannoury:1938:EOT

[228] G. Mannoury. Enige opmerkingen over taalgradatie. (Dutch) [Any comments about language degradation]. Synthese, 3(1):372–375, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00930261.

Redactie:1938:RW

[229] Redactie. Religieuse wijsbegeerte. (Dutch) [Religious philosophy]. Synthese, 3(1):376, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930262.

Hartog:1938:WBV

[230] Prof. Dr. A. H. De Hartog. De wijsgeerige beginselen van het Christendom. (Dutch) [Philosophical principles of Christianity]. Synthese, 3 (1):368–378, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930263.

deHartog:1938:PPD

[231] Prof. Dr. A. H. de Hartog. Les principes philosophiques du Christianisme. (French) [Philosophical principles of Christianity]. Synthese, 3(1): 396, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930264.

Spranger:1938:PT

[232] Eduard Spranger. Zum problem der teleologie. (German) [To the problem of teleology]. Synthese, 3(1):397–410, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930265.

Anonymous:1938:BTP

[233] Anonymous. Bijdrage tot het probleem der teleologie. (Dutch) [Contribution to the problem of teleology]. Synthese, 3(1):411–413, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930266.

Holstijn:1938:HUPc

[234] Dr. A. J. Westerman Holstijn. Hoofdstukken uit de psychoanalytische criminologie. (Dutch) [Chapters from psychoanalytic criminology]. Synthese, 3(1):414–424, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930267.

Buys:1938:MVA

[235] Dr. W. R. van Brakell Buys. De manifestatie van het absolute. (Dutch) [Absolute manifestation]. Synthese, 3(1):425–439, December 1938. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930268.

Anonymous:1938:OBG

[236] Anonymous. Ontvangen boeken en geschriften. (Dutch) [Books received and written]. Synthese, 3(1):440, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930269.

Anonymous:1938:Eb

[237] Anonymous. Errata. Synthese, 3(1):440, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930270;   http://link.
springer.com/content/pdf/10.1007/BF00930270.pdf.

Anonymous:1938:PVE

[238] Anonymous. De plaats van Eduard Spranger in de huidige psychologie. (Dutch) [The place of Eduard Spranger in current psychology]. Synthese, 3(1):441–454, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930271.

Anonymous:1938:WVE

[239] Anonymous. Werken van Eduard Spranger. (Dutch) [Works by Eduard Spranger]. Synthese, 3(1):455, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00930272.

Kruseman:1938:OS

[240] Dr. W. M. Kruseman. Organisme en samenleving. (Dutch) [Organism and society]. Synthese, 3(1):456–463, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930273.

Hartzfeld:1938:JT

[241] Mr. C. A. J. Hartzfeld. Over juridische typenleer. (Dutch) [About legal type teaching]. Synthese, 3(1):464–474, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930274.

Stammler:1938:GM

[242] Ir. C. J. J. Stamm’ler. Groep en maatschappij. (Dutch) [Group and society]. Synthese, 3(1):475–480, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00930275.

Anonymous:1938:BTa

[243] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 3(1):481–483, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930276.

Anonymous:1938:Bc

[244] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 3(1): 484, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930277.

Anonymous:1938:LC

[245] Anonymous. Leo Chestov. Synthese, 3(1):485–488, December 1938. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930278.

Anonymous:1938:HH

[246] Anonymous. A. H. De Hartog. Synthese, 3(1):489–492, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930279.

Anonymous:1938:LSV

[247] Anonymous. De lijn, die “synthese” volgt. (Dutch) [The line, which follows “synthesis”]. Synthese, 3(1):493–495, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930280.

Anonymous:1938:MVG

[248] Anonymous. De metaphysica van G. Heymans. (Dutch) [The metaphysics of G. G. Heymans]. Synthese, 3(1):496–505, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00930281.

Anonymous:1938:PDG

[249] Anonymous. Prof. Dr. G. Heymans. Synthese, 3(1):506–512, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00930282.

Daan:1938:WG

[250] Dr. Albert Daan. Wijsbegeerte en geneeskunde. (Dutch) [Philosophy and medicine]. Synthese, 3(1):513–523, December 1938. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00930283.

Anonymous:1938:Bd

[251] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 3(1):524– 527, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930284.

Anonymous:1938:BTb

[252] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 3(1):528–529, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00930285.

Anonymous:1938:FIC

[253] Anonymous. Fourth International Congress for the Unity of Science Girton College, Cambridge, England. July 14th to 19th, 1938. Synthese, 3(1): i–ii, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930286.

Anonymous:1938:USFa

[254] Anonymous. Unity of Science Forum 1938 July. Synthese, 3(1):iii– v, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930287.

Anonymous:1938:USFb

[255] Anonymous. Unity of Science Forum. Synthese, 3(1):vi, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00930288.

Anonymous:1938:USFc

[256] Anonymous. Unity of Science Forum 1938 August. Synthese, 3(1):A1– A8, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930289.

Anonymous:1938:USFd

[257] Anonymous. Unity of Science Forum 1938 September. Synthese, 3(1):A9– A12, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930290.

Neurath:1938:VHS

[258] Otto Neurath. Vereinheitlichungstendenzen hollaendischer signifiker. (German) [Unification tendencies of Dutch significance]. Synthese, 3(1): A13–A16, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930291.

Anonymous:1938:USFe

[259] Anonymous. Unity of Science Forum 1938 October. Synthese, 3(1):A17– A24, December 1938. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00930292.

Anonymous:1938:USFf

[260] Anonymous. Unity of Science Forum 1938 November. Synthese, 3(1): A25–A32, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930293.

Anonymous:1938:USFg

[261] Anonymous. Unity of Science Forum 1938 December. Synthese, 3(1): A33–A40, December 1938. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00930294.

Anonymous:1939:RVVa

[262] Anonymous. Religieuze vragen van onzen tijd. (Dutch) [Religious questions of our time]. Synthese, 4(1):3–4, December 1939. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879166.

Boerma:1939:RWD

[263] N. Westendorp Boerma. Religieus-wijsgeerig denken in onzen tijd. (Dutch) [Religious-wise thinking in our time]. Synthese, 4(1):5–8, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879167.

Anonymous:1939:MVG

[264] Anonymous. De Metaphysica van G. Heymans. (Dutch) [G. Heymans’s metaphysics]. Synthese, 4(1):9–23, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879168.

Anonymous:1939:HCD

[265] Anonymous. Hedendaagsche cultuurphilosophie. (Dutch) [Contemporary culture philosophy]. Synthese, 4(1):24, December 1939. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879169.

Breysig:1939:GMP

[266] Prof. Dr. Kurt Breysig. Die Gesamtgeschichte der Menschheit als Problem der Geschichtswissenschaft. (German) [The overall history of mankind as a problem of history science]. Synthese, 4(1):25–31, December 1939. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879170.

Groot:1939:ERI

[267] Dr. H. Groot. Emotief, rationeel en ideatief denken. (Dutch) [Emotional, rational and ideal thinking]. Synthese, 4(1):32–41, December 1939. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879171.

Busch:1939:GJV

[268] Jos F. Busch. Het gouden jaar van Bergson. (Dutch) [The golden age of Bergson]. Synthese, 4(1):42–51, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879172.

Anonymous:1939:ISS

[269] Anonymous. Internationale Signifische Studiegroep (I.S.S.). (Dutch) [International Significant Study Group (I.S.S.G.)]. Synthese, 4(1):52, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879173.

Esch:1939:TMS

[270] W. Esch. Taal, moraal en signifika als doeltreffende methode van onderzoek. (Dutch) [Language, morality and significance as effective method of research]. Synthese, 4(1):53–59, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879174.

Anonymous:1939:GIP

[271] Anonymous. Groupe international pour l’Étude de la psychologie linguistique. (French) [International group for the study of linguistic psychology]. Synthese, 4(1):59, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879175.

Gr:1939:BDB

[272] H. Gr. Boekbespreking. (Dutch) [Book review]. Synthese, 4(1):60–61, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879176.

Anonymous:1939:OBGa

[273] Anonymous. Ontvangen boeken en geschriften. (Dutch) [Books received and written]. Synthese, 4(1):61, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879177.

Anonymous:1939:BTDa

[274] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):62–64, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879178.

Boerma:1939:VND

[275] N. Westendorp Boerma. Vrijheid en noodzakelijkheid. (Dutch) [Freedom and necessity]. Synthese, 4(1):65–71, December 1939. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879179.

Jordan:1939:CCS

[276] H. J. Jordan. Causaliteit en causale structuren; de beteekenis van deze begrippen voor de wetenschap van het leven. (Dutch) [Causality and causal structures; the meaning of these concepts for the science of life]. Synthese, 4(1):72–95, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879180.

Anonymous:1939:EMO

[277] Anonymous. Ernst Mach en onze tijd. (Dutch) [Ernst Mach and our time]. Synthese, 4(1):96–105, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879181.

Anonymous:1939:UEM

[278] Anonymous. Uit: Ernst Mach, Analyse der Empfindungen. (German) [Out: Ernst Mach, analysis of sensations]. Synthese, 4(1):106–107, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879182.

Anonymous:1939:BDB

[279] Anonymous. Boekbespreking. (Dutch) [Book review]. Synthese, 4(1): 108–109, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879183.

Anonymous:1939:BTDb

[280] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):110–112, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879184.

Masson-Oursel:1939:LLB

[281] Paul Masson-Oursel. Lucien Levy-Bruhl 1857–1939. Synthese, 4(1):113– 115, December 1939. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879185.

Boerma:1939:ZBV

[282] N. Westendorp Boerma. Over de zedelijke beoordeeling van oorlog. (Dutch) [About the sudden judgment of war]. Synthese, 4(1):116– 122, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879186.

vanBemmelen:1939:WWF

[283] Dr. W. van Bemmelen. Wat is de wereld?. (Dutch) [What is the world?]. Synthese, 4(1):123–132, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879187.

Anonymous:1939:SEV

[284] Anonymous. Sociale en economische vragen. (Dutch) [Social and economic questions]. Synthese, 4(1):133–134, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879188.

Anonymous:1939:IVI

[285] Anonymous. Het Instituut Voor Internationale Industrieele Betrekkingen. (Dutch) [The Institute for International Industrial Relations]. Synthese, 4 (1):135–136, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879189.

Henderson:1939:IVK

[286] Ir. Fred Henderson. De invloeden van de krachtproductie op de economische organisatie. (Dutch) [Power production influences on economic organization]. Synthese, 4(1):137–139, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879190.

Henderson:1939:EPD

[287] Fred Henderson. The effects of power development on economic organization. Synthese, 4(1):140–157, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879191.

Daan:1939:STS

[288] Albert Daan. Strekking en Taak der Significa Algemeen Gedeelte. (Dutch) [String and Task of Significa General Section]. Synthese, 4(1):158– 169, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879192.

Daan:1939:SPG

[289] Albert Daan. Signifische probleemstudie in de geneeskunde. (Dutch) [Significant problem study in medicine]. Synthese, 4(1):170–175, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00879193.

Daan:1939:SAC

[290] Dr. Albert Daan. Signific analysis and clearing in medicine. Synthese, 4 (1):176, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879194.

Anonymous:1939:BTa

[291] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):177–179, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879195.

Anonymous:1939:UKD

[292] Anonymous. Uit het Katholieke denkleven. (Dutch) [From the Catholic thinking of life]. Synthese, 4(1):181–183, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879196.

Masson-Oursel:1939:PPF

[293] Paul Masson-Oursel. La psychologie de la poesie. (French) [The psychology of poetry]. Synthese, 4(1):184–187, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00879197.

Anonymous:1939:BTF

[294] Anonymous. Begripsvorming en taal. (Dutch) [Conceptualization and language]. Synthese, 4(1):188–201, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879198.

Anonymous:1939:FIL

[295] Anonymous. Formation d’idées et langage. (French) [Formation of ideas and language]. Synthese, 4(1):202–204, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879199.

Rautenstrauch:1939:ROA

[296] Walter Rautenstrauch. The role of organization in attaining optimum productivity. Synthese, 4(1):205–217, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879200.

Rautenstrauch:1939:FVO

[297] Prof. Walter Rautenstrauch. De functie van organiastie bij het verkrijgen van de gunstigste productiviteit. (Dutch) [Organizational function in getting the best productivity]. Synthese, 4(1):218, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00879201.

Rautenstrauch:1939:BNR

[298] Prof. Walter Rautenstrauch. Biographical note regarding Prof. Walter Rautenstrauch. Synthese, 4(1):218, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879202.

Anonymous:1939:BTb

[299] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):219–220, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879203.

Os:1939:PVW

[300] Prof. Dr. C. H. Van Os. Over de philosophie van A. N. Whitehead. (Dutch) [On the philosophy of Alfred North Whitehead]. Synthese, 4 (1):221–236, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879204.

Anonymous:1939:PW

[301] Anonymous. On the philosophy of A. N. Whitehead. Synthese, 4(1):237– 238, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879205.

Anonymous:1939:RVVb

[302] Anonymous. Religieuze vragen van onzen tijd. (Dutch) [Religious questions of our time]. Synthese, 4(1):239–243, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00879206.

Kruseman:1939:LNL

[303] Dr. W. M. Kruseman. Leven en niet-leven. (Dutch) [Life and non-life]. Synthese, 4(1):244–253, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879207.

Anonymous:1939:MF

[304] Anonymous. Mary L. Fledderus. Synthese, 4(1):254, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879208.

Fledderus:1939:OP

[305] Mary L. Fledderus. Optimum productivity. Synthese, 4(1):255–261, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879209.

Anonymous:1939:BTc

[306] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):262–263, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879210.

Anonymous:1939:EZF

[307] Anonymous. Entelechie en ziel. (Dutch) [Entelechie and soul]. Synthese, 4 (1):265, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879211.

Driesch:1939:ESG

[308] Prof. Dr. Hans Driesch. Entelechie und Seele. (German) [Entelechy and soul]. Synthese, 4(1):266–279, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879212.

Vos:1939:A

[309] Dr. H. De Vos. Autonomie. (Dutch) [Autonomy]. Synthese, 4(1):280– 291, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879213.

Anonymous:1939:P

[310] Anonymous. Plotinus. Synthese, 4(1):292, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879214.

Buys:1939:MKB

[311] Dr. W. R. Van Brakell Buys. Mensch en kosmos bij Plotinus. (Dutch) [Men and cosmos at Plotinus]. Synthese, 4(1):293–308, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879215.

Anonymous:1939:BTd

[312] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):309–310, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879216.

Anonymous:1939:HVR

[313] Anonymous. De Hoofddirectie van het Rijksmuseum te Amsterdam verzoekt ons om opname van het volgende. (Dutch) [The Main Board of the Rijksmuseum at Amsterdam requests us to record the following]. Synthese, 4(1):311, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879217.

Boerma:1939:RVV

[314] N. Westendorp Boerma. Religieuze vragen van onzen tijd. (Dutch) [Religious questions of our time]. Synthese, 4(1):313, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00879218.

Dooyeweerd:1939:TCV

[315] Prof. Dr. H. Dooyeweerd. De transcendentale critiek van het wijsgeerig denken. (Dutch) [The transcendental criticism of philosophical thinking]. Synthese, 4(1):314–339, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879219.

Waismann:1939:NPPa

[316] Dr. Friedrich Waismann. Von der Natur Eines Philosophischen Problems. (German) [The nature of a philosophical problem]. Synthese, 4(1):340– 350, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879220.

Holstijn:1939:HUP

[317] Dr. A. J. Westerman Holstijn. Hoofdstukken uit de psychoanalytische criminologie. (Dutch) [Chapters from psychoanalytic criminology]. Synthese, 4(1):351–364, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879221.

Anonymous:1939:BTe

[318] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):365–366, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879222.

Boerma:1939:AHF

[319] N. Westendorp Boerma. Autonomie en heteronomie. (Dutch) [Autonomy and heteronomy]. Synthese, 4(1):369–373, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00879223.

Anonymous:1939:CWG

[320] Anonymous. De Crisis der Wetenschap. (Dutch) [The crisis of science]. Synthese, 4(1):374–375, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879224.

Rothenberg:1939:KWL

[321] Dr. Benno Rothenberg. Zur Krise der Wissenschaft Und Ueber den Logistischen Neo–Positivismus. (German) [The crisis of science and logistic neo-positivism]. Synthese, 4(1):376–383, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879225.

Rothenberg:1939:CWG

[322] Dr. Benno Rothenberg. De Crisis der Wetenschap. (Dutch) [The Crisis of Science]. Synthese, 4(1):384–385, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879226.

Varia:1939:NMFa

[323] Signifische Varia. Nu en morgen. (Dutch) [Now tomorrow]. Synthese, 4 (1):386–387, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879227.

Mannoury:1939:NMFa

[324] Prof. G. Mannoury. Nu en morgen. (Dutch) [Now tomorrow]. Synthese, 4 (1):388–394, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879228.

Waismann:1939:NPPb

[325] Dr. Friedrich Waismann. Von der Natur Eines Philosophischen Problems. (German) [The nature of a philosophical problem]. Synthese, 4(1):395– 406, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879229.

Anonymous:1939:BTg

[326] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):407–408, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879230.

Anonymous:1939:OBGb

[327] Anonymous. Ontvangen boeken en geschriften. (Dutch) [Books received and written]. Synthese, 4(1):408–409, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879231.

Godefroy:1939:ESZ

[328] Dr. J. C. L. Godefroy. Eerste Signifische Zomerconferentie. (Dutch) [First Significant Summer Conference]. Synthese, 4(1):409–412, December 1939.

CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879232.

Anonymous:1939:M

[329] Anonymous. Massapsychologie. (Dutch) [Mass psychology]. Synthese, 4 (1):413–415, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879233.

Baschwitz:1939:MMG

[330] Dr. Kurt Baschwitz. Der Mensch in der Masse. (German) [Man in the mass]. Synthese, 4(1):416–424, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879234.

Baschwitz:1939:MMF

[331] Dr. Kurt Baschwitz. De mensch in de massa. (Dutch) [The man in the mass]. Synthese, 4(1):425, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879235.

Boerma:1939:BVT

[332] N. Westendorp Boerma. Het belang van de taalkundige psychologie. (Dutch) [The importance of linguistic psychology]. Synthese, 4(1):426– 429, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879236.

OrtegayGasset:1939:VVL

[333] Prof. Jose Ortega y Gasset. Vragen van levens- en wereldbeschouwing. (Dutch) [Questions of life and worldview]. Synthese, 4(1):430, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00879237.

OrtegayGasset:1939:MTD

[334] Prof. Jose Ortega y Gasset. De mensch en de techniek. (Dutch) [Man and technique]. Synthese, 4(1):430–433, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879238.

Varia:1939:NMFb

[335] Signifische Varia and Prof. G. Mannoury. Nu en morgen. (Dutch) [Now tomorrow]. Synthese, 4(1):434–449, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879239.

Anonymous:1939:BTh

[336] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):450–454, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879240.

Anonymous:1939:HDW

[337] Anonymous. Hans Driesch als wijsgeer. (Dutch) [Hans Driesch as philosopher]. Synthese, 4(1):455–466, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879241.

Driesch:1939:IS

[338] Prof. Dr. Hans Driesch. II. Das “Sein”. (German) [II. The being]. Synthese, 4(1):467–482, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879242.

Anonymous:1939:C

[339] Anonymous. Causerie: Gehouden door Prof. G. Mannoury op de Eerste Signifische Zomerconferentie te Naarden op 15 Juli 1939. (Dutch) [causerie: Held by Prof. G. Mannoury at the First Significant Summer Conference in Naarden on July 15, 1939]. Synthese, 4(1):483– 485, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879243.

Jordan:1939:SBG

[340] Herm J. Jordan. Signifik und Biologie. (German) [Significance and biology]. Synthese, 4(1):486–503, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879244.

Jordan:1939:SBF

[341] Prof. Dr. H. J. Jordan. Significa in de biologie. (Dutch) [Significance in biology]. Synthese, 4(1):504–508, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879245.

Stoop:1939:NJV

[342] A. Stoop, Jr. and J. Mannoury. In het “Nederlandsch Juristenblad” van 28 October 1939 verscheen de volgende oproep. (Dutch) [In the “Dutch Juridical Journal” of 28 October 1939 the following call appeared]. Synthese, 4(1):509–510, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879246.

Berdjajew:1939:DPV

[343] Prof. N. Berdjajew. Dostojewsky en het probleem der vrijheid. (Dutch) [Dostoevsky and the problem of freedom]. Synthese, 4(1):511–526, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879247.

Berdjajew:1939:DPF

[344] Prof. N. Berdjajew. Dostojewsky und das Problem der Freiheit. (German) [Dostoevsky and the problem of freedom]. Synthese, 4(1):527– 528, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879248.

Anonymous:1939:BTi

[345] Anonymous. Buitenlandsche tijdschriften. (Dutch) [Foreign magazines]. Synthese, 4(1):529–533, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879249.

Anonymous:1939:GVS

[346] Anonymous. De Gang van “Synthese”. (Dutch) [The operation of “Synthese]. Synthese, 4(1):535–539, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879250.

Boerma:1939:EWJ

[347] N. Westendorp Boerma. De ethische wijsgeer I. J. de Bussy. (Dutch) [The ethical sage I. J. de Bussy]. Synthese, 4(1):540–543, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879251.

deBussy:1939:EAF

[348] I. J. de Bussy. Egoisme en altruisme. (Dutch) [Egoism and altruism]. Synthese, 4(1):544–553, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879252.

Friedman:1939:BL

[349] Dr. D. Friedman. Het buddhisme als levenspad. (Dutch) [Buddhism as life path]. Synthese, 4(1):554–562, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879253.

Mannoury:1939:NMFb

[350] G. Mannoury. Nu en morgen. (Dutch) [Now tomorrow]. Synthese, 4 (1):563–570, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879254.

Meyer:1939:NLG

[351] H. Meyer. Naturwissenschaft und Logik. (German) [Natural science and logic]. Synthese, 4(1):599–600, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00879255.

OrtegayGasset:1939:CID

[352] Prof. Jose Ortega y Gasset. Collectivisme en individualisme. (Dutch) [Collectivism and individualism]. Synthese, 4(1):601–610, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879256.

OrtegayGasset:1939:OKG

[353] Prof. Jose Ortega y Gasset. Die Organisation des Kollektivlebens. (German) [The organization of collective life]. Synthese, 4(1):611–612, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879257.

Anonymous:1939:USFa

[354] Anonymous. Unity of Science Forum 1939 January. Synthese, 4(1):A41– A43, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879258.

Grelling:1939:LSG

[355] Kurt Grelling. Zur Logik der Sollsaetze. (German) [On the logic of the set points]. Synthese, 4(1):A44–A47, December 1939. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00879259.

Anonymous:1939:BA

[356] Anonymous. Books and articles. Synthese, 4(1):A47–A48, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00879260.

Anonymous:1939:USFb

[357] Anonymous. Unity of Science Forum 1939 February. Synthese, 4(1): A49–A56, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879261.

Anonymous:1939:USFc

[358] Anonymous. Unity of Science Forum 1939 March. Synthese, 4(1):A57– A60, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879262.

Neurath:1939:IEE

[359] Otto Neurath. Die Internationale Enzyklopaedie der Einheitswissenschaft. (German) [The International Encyclopedia of Unity Science]. Synthese, 4(1):A61, December 1939. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00879263.

Kraft:1939:OLE

[360] J. Kraft. Orthodoxer und Liberaler Empirismus. (German) [Orthodox and liberal empiricism]. Synthese, 4(1):A61–A64, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00879264.

Anonymous:1939:FIC

[361] Anonymous. Fifth International Congress for the Unity of Science. Synthese, 4(1):A64, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00879265.

Anonymous:1939:USFd

[362] Anonymous. Unity of Science Forum 1939 April. Synthese, 4(1):A65– A72, December 1939. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00879266.